OpenStudy (supersquidsisters):

help? i keep getting the wrong answer? What is the 32nd term of the arithmetic sequence where a1 = –32 and a9 = –120? (1 point) –384 –373 –362 –351

OpenStudy (faiqraees):

Can you show us your working?

OpenStudy (supersquidsisters):

@FaiqRaees im using an = a1 + (n - 1)d a32 = -32 + (-120 - 1)d but i dont know what to put in place of d

OpenStudy (supersquidsisters):

i actually know why i was getting the wrong answer at first but i cant get past here

OpenStudy (faiqraees):

Okay so you are mistaking things N is not equal to a term in the series. It describes the position of the number in a series D is the difference between term a1 implies first term, a2 the second term and so on So your first task is to find the difference. try finding it out by plugging a9 and a1 in the formula

OpenStudy (faiqraees):

aN is the number of your choice from the sequence.

OpenStudy (supersquidsisters):

@FaiqRaees in this case shouldnt it be 32 since thats the number youre trying to find in the sequence?

OpenStudy (faiqraees):

Yes true, but if we directly start with a32, we have to account for the unknown value of 'd'. So first we have to figure out the value of 'd' : difference

OpenStudy (supersquidsisters):

@FaiqRaees aaaah i see what youre sayin now. so what do you need to find the difference between??

OpenStudy (faiqraees):

\[\large\rm a_9=a_1(n-1)d\ \]

OpenStudy (supersquidsisters):

@FaiqRaees so would i substitute into that equation? so a9 = -32 (9 - 1)d ???

OpenStudy (faiqraees):

You have the value of a9 right? Use that and find out the difference

OpenStudy (supersquidsisters):

@FaiqRaees oh oops so -120 = -32 (-120 - 1)d ??? idk if thats right though because every time i work it out i get some long decimal

OpenStudy (faiqraees):

What is the value of 'n' here?

OpenStudy (supersquidsisters):

@FaiqRaees n would be a9 wouldn't it? since you said it wasnt the number were looking for

OpenStudy (faiqraees):

n is the term number

OpenStudy (faiqraees):

Still confused?

OpenStudy (phi):

you use \[ a_n = a_1 + (n - 1)d \] you have to be meticulous (careful). they give you \(a_9\) which is associated with n=9 (the n comes from the 9 in a9) so using a1 = –32 and a9 = –120 we use n=9, and -120= -32+(9-1)*d can you "solve for d" ?

OpenStudy (supersquidsisters):

@phi @FaiqRaees oh ok that makes sense um so d=5?

OpenStudy (faiqraees):

No d is not 5. Can you show the working for d?

OpenStudy (supersquidsisters):

@FaiqRaees ah dang sorry im so bad at math omg its so embarrassing im always screwing things up ughgghhg -120 = -32 + (9 - 1)d -120 = -32 + (8)d -120 = -24d 5 = d im sorry

OpenStudy (phi):

-120= -32+(9-1)*d first simplify 9-1 to 8 -120 = -32+8d now add +32 to both sides. then simplify both sides.

OpenStudy (phi):

your work looks good except you -32+8d is not -24d for the time being, just believe that, and instead, add +32 to both sides

OpenStudy (supersquidsisters):

@phi so -120 = -32 + 8d -88 = 8d ???

OpenStudy (phi):

yes. now divide both sides by 8 \[ \frac{-88}{8}= \frac{8}{8}d \]

OpenStudy (supersquidsisters):

@phi d = -11??

OpenStudy (faiqraees):

Correct

OpenStudy (phi):

yes now the formula is \[ a_n= -32+ (n-1)\cdot -11 \]

OpenStudy (faiqraees):

Now you have a1 and difference. Try finding out the value of a32

OpenStudy (phi):

can you find a32 ? (i.e. use n=32)

OpenStudy (supersquidsisters):

@phi @FaiqRaees so would a32 = -374 ??

OpenStudy (faiqraees):

-373, do you mean?

OpenStudy (phi):

did you do a32= -32 + -11*(32-1) = -32 + -11*31 = -32 - 341 = -373

OpenStudy (phi):

btw, if you have time, see https://www.khanacademy.org/math/algebra-basics/core-algebra-linear-equations-inequalities/core-algebra-solving-basic-equations/v/equations-2 for how to do the algebra. Your life will be easier if you learn how to do that.

OpenStudy (supersquidsisters):

@phi oh i got 342 instead of 341 for some reason :/

OpenStudy (phi):

did you do -11*32 by hand or with a calculator?

OpenStudy (phi):

* I mean -11*31

OpenStudy (supersquidsisters):

@phi by a calculator but i typed it in wrong i see where i messed up oops